The rate of economic growth per capita in France from 1996 to 2000 was 1.9% per year, while in Korea over the same period it was 4.2%. Per capita real GDP was $28,900 in France in 2003, and $12,700 in Korea. Assume the growth rates for each country remain the same.
Compute the doubling time for France’s per capita real GDP.
Compute the doubling time for Korea’s per capita real GDP.
What will France’s per capita real GDP be in 2045?
What will Korea’s per capita real GDP be in 2045?
1
Expert's answer
2015-03-13T09:43:11-0400
The rate of economic growth per capita in France from 1996 to 2000 was 1.9% per year, while in Korea over the same period it was 4.2%. Per capita real GDP was $28,900 in France in 2003, and $12,700 in Korea.
We use the rule of 70.
The doubling time for France’s per capita real GDP is 70/1.9 = 36.8 years, so France's GDP will double in 2040.
The doubling time for Korea’s per capita real GDP is 70/4.2 = 16.7 years, so France's GDP will double in 2020.
France’s per capita real GDP in 2045 will be 28,900*1.019^42 = $63,710.88
Korea’s per capita real GDP in 2045 will be 12,700*1.042^42 = $71490.43